An ordinary (fair) die is a cube with the numbers 1 through 6 on the sides (represented by painted spots). Imagine that such a die is rolled twice in succession and that the face values of the two rolls are added together. This sum is recorded as the outcome of a single trial of a random experiment.
Compute the probability of each of the following events:
Event : The sum is greater than 5 .
Event : The sum is divisible by 6.
Write your answers as exact fractions.

Answer by edjones(8007)

sum>5
2:1/36
3:2/36
4:3/36
5:4/36
--------add
==10/36
1 - 10/36 = 26/36 = 13/18
.
sum divisible by 6
6:5/36
12:1/36
--------add
==6/36=1/6
.
Ed